# Efficient and Scalable Density Functional Theory Hamiltonian Prediction through Adaptive Sparsity

### v1

@@ Appendix B Additional Experimental Results
B.1 Ablation Study for Three-phase Sparsity Scheduler
We evaluated the impact of the three-phase sparsity scheduler on model performance. We trained the SPHNet under several settings: without the random stage, without the adaptive stage, and without the fixed stage. Each configuration was trained five times to assess overall performance. As shown in Table 5, models trained without the random stage tended to converge to localized optimal points, resulting in suboptimal performance. The models trained without an adaptive stage exhibited more randomness in performance, achieving significantly lower accuracy compared to the SPHNet with a complete scheduler. For the model trained without the fixed stage, although it did not show a substantial difference in accuracy metrics, the speedup ratio dropped to 5.45 due to the additional computational overhead from the sparse gate and the non-static graph implementation of tensor products following the Sparse Pair Gate, as opposed to standard implementations in the e3nn library.
Table 5: Ablation Experimental results on Three-phase Sparsity Scheduler.
Model Random Adaptive Fixed H $\downarrow$ Memory $\downarrow$ Speed $\uparrow$ Speedup $\uparrow$
Stage Stage Stage [$10^{-6}E_{h}$] [GB/Sample] [Sample/Sec] Ratio
SPHNet 97.31 $\pm$ 0.52 5.62 3.12 7.09x
SPHNet 112.68 $\pm$ 10.75 5.62 3.12 7.09x
SPHNet 122.79 $\pm$ 19.02 5.62 3.12 7.09x
SPHNet 97.11 $\pm$ 1.33 5.62 3.12 5.45x
B.2 Ablation Study for Sparse Gates and Interaction blocks
Table 6: Ablation Experimental results on PubChemQH dataset.
Model Sparse Pair Sparse TP Vectorial Node Spherical Node H $\downarrow$ Memory $\downarrow$ Speed $\uparrow$ Speedup $\uparrow$
Gate Gate Interaction block Interaction block [$10^{-6}E_{h}$] [GB/Sample] [Sample/Sec] Ratio
SPHNet 4 2 97.31 5.62 3.12 7.09x
SPHNet 4 2 94.31 8.04 1.75 3.98x
SPHNet 4 2 87.70 6.98 1.20 2.73x
SPHNet 4 2 86.35 10.91 0.76 1.73x
SPHNet 0 5 97.08 8.47 1.08 2.45x
To better evaluate the effects of the Sparse Pair Gate and Sparse Tensor Product Gate on the overall performance, we also trained the SPHNet on three different settings, including SPHNet without Sparse Pair Gate, SPHNet without Sparse Tensor Product Gate, and SPHNet without both Gate. As shown in Table 6, the Sparse Pair Gate could bring 78% speedup and 30% GPU memory reduction to the model with little loss of model accuracy. The Sparse Tensor Product Gate in our setting (remove 70% combinations in the tensor product) would lose about 10% accuracy, however, it could bring 160% speedup, which was enough for us to ignore the loss it caused. It’s noteworthy that when the computational resource is higher, the user can set a lower sparsity rate to improve the accuracy (details are presented in section 5.4).
We then evaluated the impact of the Vectorial Node Interaction block on model performance. As shown in Table 6, we replaced the Vectorial Node Interaction block with three Spherical Node Interaction blocks. The results indicated that the accuracy of the Hamiltonian MAE remained comparable to that of the original SPHNet, however, performance in terms of memory consumption and speed declined significantly. This suggests that, despite the absence of the tensor product operation, the Vectorial Node Interaction block can effectively extracts atomic information from the input data with much less memory consumption and much faster training speed.
B.3 Selected Combination Set Details
@@
B.4 GPU Memory when Scaling with Increasing Size
@@
B.5 Single Tensor Product Time Scaling with Sparsity Rate
@@ D.5.2 Non-Diagonal Block
For the non-diagonal pairs, the module first uses Sparse Pair Gate to select the most valuable pair $\{\mathbf{x^{\prime}}_{i},\mathbf{x^{\prime}}_{j}\}$ and their tensor product weight $\mathbf{w}_{ij}$, as defined in Equation 9. The tensor product weight is then multiplied with the RBF of $\vec{r^{\prime}_{ij}}$ and gets the final tensor product weight as defined in Equation 12. Then, the module calculates the tensor product of $\mathbf{x^{\prime}}_{i}$ and $\mathbf{x^{\prime}}_{j}$, the process is presented in Fig.9E. Noted that the combinations in the tensor product are selected by the Sparse Tensor Product Gate. The result is regarded as the desired $\mathbf{f}_{ij}$, defined in Equation 14. Last, the non-diagonal feature $\mathbf{f}_{ij}$ is outputted after normalization and sph linear the same as in the Spherical Node Interaction block.
@@ E.1 SE(3) equivariant neural network
These SE(3) equivariant neural networks greatly improve the performance of Artificial Intelligent in the field of quantum chemistry. However, the computational complexity of tensor product operation greatly reduces the efficiency of these SE(3) equivariant models. The eSCN(Passaro & Zitnick, 2023) presents an efficient method to perform SO(3) equivariant convolutions. It reduces the computational complexity by aligning node embeddings’ primary axis with edge vectors, transforming the SO(3) convolutions into mathematically equivalent SO(2) convolutions, which decreases complexity from $O(L^{6})$ to $O(L^{3})$.
@@ E.2 Hamiltonian matrix prediction
The SE(3) equivariant neural network is one of the most used models in the field of Hamiltonian matrix prediction(Unke et al., 2021; Yu et al., 2023; Gong et al., 2023; Atz et al., 2021). The PhiSNet (Unke et al., 2021) is the first method that primarily focused on Hamiltonian matrix prediction. It leverages SE(3)-equivariant operations through its whole architecture to predict molecular wavefunctions and electronic densities and can reconstruct wavefunctions with high accuracy. The main problem for PhiSNet is its inefficiency due to the large amount of tensor product operations. To solve this problem, QHNet (Yu et al., 2023) proposed a model with careful design to greatly reduce the number of tensor products and improve the efficiency of Hamiltonian prediction. The DeepH-E3 (Gong et al., 2023) is an E(3)-equivariant model that preserves Euclidean symmetry even with spin-orbit coupling. The method allows for efficient and accurate electronic structure calculations of large-scale materials by learning from small-sized DFT data, significantly reducing computational costs.
There are also other works that try to enhance the prediction ability from different aspects. DeepH (Li et al., 2022) introduces a local coordinate system defined for each edge according to its local chemical environment to handle the gauge (or rotation) covariance of the DFT Hamiltonian matrix. This allows the Hamiltonian matrix blocks to be invariant under rotation when transformed into the local coordinate system. Self-consistency training(Zhang et al., 2024) leverages the self-consistency principle of density functional theory (DFT) to train a model without requiring the labeled Hamiltonian matrice. This enhances the generalization and efficiency and reduces reliance on costly DFT calculations for supervision.

### v2

@@ Appendix B Additional Experimental Results
B.1 Ablation Study for blocks of SPHNet
We conducted an ablation study to evaluate the effect of different modules in the SPHNet architecture. Specifically, the standard SPHNet model has 4 Vectorial Node Interaction blocks, 2 Spherical Node Interaction blocks, and 2 Pair Construction blocks. We removed all the sparse gates and reduced the number of these three kinds of modules to 1, respectively, and observed the model performance. As shown in the table below, we found that both the Vectorial Node Interaction block and the Spherical Node Interaction block significantly affect the model performance, indicating that the design of architectures with progressively increased irreps orders has an important positive impact on the models. Interestingly, we found that removing one Pair Construction block would not strongly affect the model accuracy, suggesting that there is actually room to further speed up the model. We will explore this further in our future work.
Table 8: The effect of blocks’ numbers in SPHNet on the PubChemQH dataset.
Model Sparse Pair Sparse TP Vectorial Node Spherical Node Pair Construction H $\downarrow$
Gate Gate Interaction block Interaction block block [$10^{-6}E_{h}$]
SPHNet 4 2 2 86.35
SPHNet 1 2 2 96.01
SPHNet 4 1 2 97.35
SPHNet 4 2 1 89.17
B.2 The Effect of Sparity Rate on the Computational Cost
We analyze the effect of sparsity rate on the model performance in Section 5.4. Here we put the complete results of training speed and GPU memory in the PubChemQH dataset under every sparse rate. The results showed that the computational cost reduced linearly as the sparsity increased, which was in line with our expectations.
Table 9: The effect of sparsity rate v.s. computational cost on PubChemQH dataset
Sparsity H $[10^{-6}E_{h}]$ Speed [Sample/Sec] Memory [GB/Sample]
0 78.48 0.78 17.02
10 79.90 0.96 12.49
20 81.29 1.05 11.21
30 85.33 1.24 10.93
40 86.65 1.50 6.84
50 90.52 1.86 6.22
60 91.90 2.31 5.12
70 97.31 3.12 5.62
80 158.70 3.92 4.68
90 183.52 5.13 3.38
B.3 Single Tensor Product Time Scaling with Sparsity Rate
B.4 Selected Combination Set Details
@@
B.5 GPU Memory when Scaling with Increasing Size
@@ D.5.2 Non-Diagonal Block
For the non-diagonal pairs, the module first uses Sparse Pair Gate to select the most valuable pair $\{\mathbf{x^{\prime}}_{i},\mathbf{x^{\prime}}_{j}\}$ and their tensor product weight $\mathbf{w}_{ij}$, as defined in Equation 9. The tensor product weight is then multiplied with the RBF of $\vec{r^{\prime}_{ij}}$ and gets the final tensor product weight as defined in Equation 12. Then, the module calculates the tensor product of $\mathbf{x^{\prime}}_{i}$ and $\mathbf{x^{\prime}}_{j}$, the process is presented in Fig.9E. Note that the combinations in the tensor product are selected by the Sparse Tensor Product Gate. The result is regarded as the desired $\mathbf{f}_{ij}$, defined in Equation 14. Last, the non-diagonal feature $\mathbf{f}_{ij}$ is outputted after normalization and sph linear the same as in the Spherical Node Interaction block.
@@ D.6 Expansion Block
D.7 Computational Overhead of the Sparse Gates and Scheduler
Compared to the tensor product operation, Sparse Gate and the three-phase learning schedule only introduce minimal computational overhead and have little impact on the overall speed. We would like to discuss this part of computational cost here.
In the three-phase sparsity scheduler, for a given unsparsified set $U$, the additional computational overhead in the first phase has a complexity of $\mathcal{O}(|U|)$, contributed by the RANDOM(·) operation. The second phase has a computational overhead of $\mathcal{O}(|U|\log{|U|})$, arising from the TOP(·) operation. Since we fix the learnable weight matrix and the selected elements, there is no additional computational overhead in the third phase. For detailed information, please refer to Equation 3.
For the sparse TP gate, the computational overhead comes from the element-wise multiplication of two weight vectors (Equation 5), so its complexity is $\mathcal{O}(|U_{c}|)=\mathcal{O}(L^{3})$.
For the sparse pair gate, the additional computational overhead mainly comes from the linear layer $F_{p}(\cdot)$ in Equation 7, with its time complexity being the square of its hidden dimension. Other operations, including the inner product (Equation 6) and the weight calculation (Equation 9), are necessary operations in our framework, even without the sparse pairwise gate.
@@ E.1 SE(3) Equivariant Neural Network
These SE(3) equivariant neural networks greatly improve the performance of Artificial Intelligent in the field of quantum chemistry. However, the computational complexity of tensor product operation greatly reduces the efficiency of these SE(3) equivariant models. The eSCN (Passaro & Zitnick, 2023) presents an efficient method to perform SO(3) equivariant convolutions. It reduces the computational complexity by aligning node embeddings’ primary axis with edge vectors, transforming the SO(3) convolutions into mathematically equivalent SO(2) convolutions, which decreases complexity from $O(L^{6})$ to $O(L^{3})$. E2Former (Li et al., 2025a) introduces an equivariant and efficient transformer architecture that incorporates the Wigner 6j convolution (Wigner 6j Conv). By shifting the computational burden from edges to nodes, the Wigner 6j Conv reduces the complexity from $O(|E|)$ to $O(|V|)$ while preserving both the model’s expressive power and rotational equivariance.
@@ E.2 Hamiltonian Matrix Prediction
The SE(3) equivariant neural network is one of the most used models in the field of Hamiltonian matrix prediction (Unke et al., 2021; Yu et al., 2023; Gong et al., 2023; Atz et al., 2021). The PhiSNet (Unke et al., 2021) is the first method that primarily focuses on Hamiltonian matrix prediction. It leverages SE(3)-equivariant operations through its whole architecture to predict molecular wavefunctions and electronic densities, and can reconstruct wavefunctions with high accuracy. The main problem for PhiSNet is its inefficiency due to the large amount of tensor product operations. To solve this problem, QHNet (Yu et al., 2023) proposed a model with careful design to greatly reduce the number of tensor products and improve the efficiency of Hamiltonian prediction. The DeepH-E3 (Gong et al., 2023) is an E(3)-equivariant model that preserves Euclidean symmetry even with spin-orbit coupling. The method allows for efficient and accurate electronic structure calculations of large-scale materials by learning from small-sized DFT data, significantly reducing computational costs.
There are also other works that try to enhance the prediction ability from different aspects. DeepH (Li et al., 2022) introduces a local coordinate system defined for each edge according to its local chemical environment to handle the gauge (or rotation) covariance of the DFT Hamiltonian matrix. This allows the Hamiltonian matrix blocks to be invariant under rotation when transformed into the local coordinate system. Self-consistency training (Zhang et al., 2024) leverages the self-consistency principle of density functional theory (DFT) to train a model without requiring the labeled Hamiltonian matrice. This enhances the generalization and efficiency and reduces reliance on costly DFT calculations for supervision. The DEHQ (Wang et al., 2024) integrates Deep Equilibrium Models (DEQs) to predict Hamiltonians. It inherently captures the self-consistent nature of Hamiltonians, a critical aspect often overlooked by traditional machine learning methods. By employing DEQs, the model circumvents the need for iterative DFT calculations during training. WANet (Li et al., 2025b) introduces a scalable deep learning model for Hamiltonian prediction. By leveraging a novel Wavefunction Alignment Loss (WALoss), the model notably reduces total energy error derived from the predicted Hamiltonian matrix and accelerates SCF calculations with the predicted Hamiltonian matrix.